LUCIFER Technical Observing, prep for Partner Science:
UT 2010 Sep 29

Observer: DThompson
Telescope Operator: SAllanson, GBechetti
Instrument Support: TShih
Telescope Support: MMidkiff
SW support: MDeLaPena

Summary

Preparations

Details

23:40 Had to restart ReadoutService to get archive information to take.
23:45 Restarted TelescopeService as TCS info was not being read.
00:00 Closed-dome pupil image saturates with lights on. Must wait.
00:35 LBTtools updated to include Sandbox/rangebal task
01:07 Dome is open. LUCI#0002-0004 are with dome closed or partly open but M3 is out of position.
01:15 LUCI#0005 at zenith. Crane is partially in beam? Claim is no after checking. Pupil is way off.
  • pupil0005.png

Images 0008 (rot=341) and 0009 (rot=161) taken at maximum M3 offset. Pupil moves a lot. Note that M3 positions are recorded in image headers.

02:21 Preset to ACT0234 to correct pointing. Did a ptreset before the slew. Two equal-brightness stars in field.
02:23 Preset to ACT0198. Pointing corrected by ~+13 in IE and ~-11 in CA.

02:26 Preset to BS9178 ACTIVE to collimate. Initial collimation not bad (blobby but <1" FWHM on guider. First WFS cycle 1612nm rms).

02:31 Preset to BS9178 using the on-axis star for the WFS. PA=180 and ROT=173 last WFS image from this set is 0011. Preset to PA=0, first wfs image here is 0012. Sources in subaps moved 0.3 pix in X and 1.4 pix in Y on the WFS chip.

02:51 Take on-sky image with Luci - #0015.

03:36 Preset to BS9178 with WFS on axis:
  • PA=0, ROT=-25 is #0018
  • PA=90, ROT=65 is #0019
  • PA=180, ROT=155 is #0020
  • PA=270, ROT=244 is #0021
  • PA=0, ROT=333 is #0022
03:52 Flexure compensation on
Telescope Config LUCI# Cold Stop Ctr M2refl Ctr
PA=0, ROT=332 #0023 938.0 1075.0 838.0 1093.0
PA=270, ROT=242 #0024 938.0 1077.0 844.5 1103.0
PA=180, ROT=151 #0025 932.0 1046.0 844.5 1054.0
PA=90, ROT=61 #0026 932.5 1044.0 837.5 1043.5
PA=0, ROT=-29 #0027 935.0 1074.5 833.0 1092.5

04:06 ACT0398 to correct pointing, #0028

04:10 BS9177 for field measurements

6:00 This turned into a bit of a debugging session for IDL, meanwhile Dave T has worked out some M3 changes to try that match nicely to my expectations for M3 so we are going to try adjusting M3 to minimize telescope alignment induced precession of the M2 image in the Lucifer pupil image, and once converged then take some field data.

06:13 Moving to BS9188.

Telescope Config LUCI# Cold Stop Ctr M2refl Ctr
PA=180, ROT=370 #0030    
PA=90, ROT=285 #0031    
PA=0, ROT=196 #0032    

M3TIP changed from 75 to -17

PA=0 ROT=201 LUCI#33
PA=180 ROT=23 LUCI#34

M3TILT changed from 0 to 153

PA=180 ROT=27 L#0035
PA=0 ROT=210 L#0036

PA=0 ROT=211 L#0037
PA=180 ROT=38 L#0038

PA=150 L#0039
PA=-30 L#0040

Retake the 4 rotation positions, correct for cold-stop wander PA=180 ROT=44 L#0041
PA=270 ROT=136 L#0042
PA=0 ROT=227 L#0043
PA=90 ROT=319 L#0044

M3TIP set to -17, M3TILT set to -75 (total collimation values). With this M3 position, we measured for four rotator positions the centroid of the M2 reflection and the center of the cold stop in LUCIFER. Assume the cold stop moves from internal LUCIFER flexures that also affect the M2 reflection. Measurements in table taken at EL~75.

Values in pixels
Luci# PA ROT Ctr of M2img Ctr of CStop dX dY
0051 100 359.8 848.5 1091.5 930.5 1075.5 -82.0 16.0
0052 10 270.2 839.5 1096.5 936.5 1077.5 -97.0 19.0
0053 280 180.5 835.0 1065.0 931.0 1045.0 -96.0 20.0
0054 190 90.9 844.5 1060.0 929.0 1045.0 -84.5 15.0

  • 51-53 difference 14.0 -4.0 (swingarm vertical)
  • 52-54 difference -12.5 4.0 (swingarm horizontal)

For 51-53 you need to move the M2 reflection to the left from the swingarm (+M3TIP) to align, plus a little -M3TILT.

For 52-54 you mostly need to move the M2 reflection away from the swingarm (-M3TILT) to align, plus a little -M3TIP.

7:40 going to guide star BS 9107 to take field data. M3 has been moved to absolute position -632" Tip and -95" Tilt.

Guide Star wfscimage000*
0 3-4
18 5-7
0 8
70 9-11
0 12
96 13-15
0 16
97 17-19
0 20
64 21-23
0 24
11 25-27
0 28
40 29-31
0 32
82 33-35
0 36
| 91 | 37-39 | Pupil centre 1 row low here, row 6?
0 40
71 41-43
0 44
72 45-47
0 48
29 49-51
0 52-57

something weird going on here, not converging well for no apparent reason and pupil centre seems to be in wrong column (8). Anyway, second round of field collect starting now... all for one big data set.
0 58
18 59-61
0 62
70 63-65
0 66
96 67-69
0 70-71
97 72-74
0 75-76
| 64 | 77-79 | slightly dubious data
0 80
| 11 | 81-83 | dome seeing? again slightly dubious data. turns out vent doors are shut, so probably there is dome seeing here. | 0 | 84 | partially opening vent doors.
40 85-87
0 88-89
82 90-92
0 93-94

9:30 field data collection complete, over to Dave.

09:40 Check accuracy of dithering on slit. Frames 67 (on-axis), 68-89 at +/- 120 arcsec in REL DETXY Y offsets, 10 cycles. All are Br_gam+clear with 6x10.0s exposures saved integrated, taken with flexure compensation ON to simulate spectroscopic observations. Point sources near top of LUCIFER image are rounder, those at the bottom are clearly elongated. Andrew suggests observing same field with 4 guide stars just off the field of view at widely spaced position angles.

10:15 L#0090 BS9107 gs=47 PA=0 [Offset: BS->GS47 = 212.11 arcsec (87.19 W 193.36 N)] PCS_XY coords: 52.311 116.018
10:19 L#0091 BS9107 gs=68 PA=180 [Offset: BS->GS68 = 178.25 arcsec (24.73 E 176.53 S)] PCS_XY coords: 14.848 105.914
10:24 L#0092 BS9107 gs=91 PA=90 [Offset: BS->GS91 = 188.41 arcsec (188.27 E 7.32 S)] PCS_XY coords: -4.399 112.960
10:29 L#0093 BS9107 gs=29 PA=270 [Offset: BS->GS27 = 230.44 arcsec (225.94 W 45.31 S)] PCS_XY coords: 27.187 135.565
PCS_XY coords from pointto's approximation. Guider showing ~0.5" or better, round, and quite stable throughout.

10:41 L#0094 Optic Sieve image (same settings as above), too faint
10:43 L#0095 broadband K Optic Sieve

10:46 L#0096 Back to off-axis star in LUCIFER catalog [BS->GS 115.29 E 263.14 N 287.29 23.66] Initial images of guide star fairly out of focus. WFS not collimating very well now. Cleared all active optics and forces. Had to reverse the focus offset put in earlier. L#0096 taken with wfs starting at ~500. Source on guide camera getting trefoil again. Definitely something odd as we were collimating on this source throughout the whole dither sequence above (or were we?). Last WFS image is 298.

Change back to gs=47 (only 212 arcsec off axis) and it is collimating but still having some trouble and not collimating very fast given the 0.5" or better seeing. At UT11:04 Steve opened the vent doors 1/2. Dome seeing? As the shutters were opening, the wfs quickly got to ~100nm rms and the guider FWHM dropped significantly. Source of dome seeing should be investigated if it is this bad!

11:05 L#0097 BS9107 gs=47 PA=0

11:10 L#0098 BS9107 gs=70 PA=0 (258" off axis at PA=7.8) Elongated sources at bottom, looks round at top
11:15 Closing left vent door to 1/2 open as wind speed is rising slightly.
11:19 L#0099 BS9107 gs=70 PA=300 sources at bottom have tails, at top they look OK
11:24 L#0100 BS9107 gs=70 PA=60, L#0101 same settings, sources still have tails. WFS looks good (low rms, round guide star).
11:28 L#0102 BS9107 gs=70 PA=300 Images round across field. Guide cam showing images as small as 0.29"!
11:33 L#0103 BS9107 gs=70 PA=0 Tails at bottom of image, as before at PA=0
11:38 L#0104 BS9107 gs=70 PA=60 All sources have triangular shape.
11:43 L#0105 BS9107 gs=70 PA=300

Check collimation model
  • Started collimated at EL~71d
  • Preset to source at EL~24 in TRACK mode, looks good! L#0106 (1x10.0s, Br_gam)
  • Preset to source at EL~43, looks good. L#0107
  • Preset to source at EL~69, looks OK (elongated, but still <5 pixels). L#0108
  • preset to source at EL~80, looks OK (same as previous) L#0109

Check pointing model
  • Collimate at EL~64
  • Started pointing log 20100929.ptlog

WT10_370 0110
ACT0406 0111
ACT0400 0112
ACT0342 0113
ACT0276 0114
ACT0191 0115
ACT0248 0116
ACT0093 0117
WT10_163 0118
WT10_269 0119
WT10_220 0120
WT10_322 0121

Turned over to Walter and Nancy so they can look into the pupil issue.

-- DavidThompson - 23 Sep 2010

  • pupil0006.png:
    pupil0006.png

  • pupil0007.png:
    pupil0007.png

  • pupil0008.png:
    pupil0008.png
I Attachment Action Size Date Who Comment
pupil0005.pngpng pupil0005.png manage 95 K 29 Sep 2010 - 01:23 DavidThompson  
pupil0006.pngpng pupil0006.png manage 72 K 29 Sep 2010 - 01:33 DavidThompson  
pupil0007.pngpng pupil0007.png manage 90 K 29 Sep 2010 - 01:40 DavidThompson  
pupil0008.pngpng pupil0008.png manage 75 K 29 Sep 2010 - 01:40 DavidThompson  
Topic revision: r10 - 29 Sep 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback